### v1

@@ Abstract
The study of undecidability in problems arising from physics has experienced a renewed interest, mainly in connection with quantum information problems. The goal of this review is to survey this recent development. After a historical introduction, we first explain the necessary results about undecidability in mathematics and computer science. Then we briefly review the first results about undecidability in physics which emerged mostly in the 80s and early 90s. Finally we focus on the most recent contributions, which we divide in two main categories: many body systems and quantum information problems.
journal: $\rightarrow$ to submit to Physics Reports.
@@ 1 Introduction
In a note about one of those attempts, C. Bennett wrote (Bennett (1990)):
Alluding to the effects on twentieth-century physics of quantum mechanics and general relativity, the cosmologist John Wheleer once remarked that the world is simpler and stranger than we can imagine. An analogous dose of strange simplicity was delivered to mathematics in the 1930s by the discoveries of Gödel, Turing and their contemporaries in the theory of proof and computation. Yet there has been surprisingly few applications of these ideas to physics. In particular, the phenomenon of ’undecidability’ has not asserted itself in a simple, uncontrived physical setting the way that deterministic chaos, a far weaker property, has.
And some years later, R. Penrose commented (Penrose (2005) p378):
It is perhaps remarkable, in view of the close relationship between mathematics and physics, that issues of such basic importance in mathematics as transfinite set theory and computability have as yet had a very limited impact on our description of the physical world. It is my own personal opinion that we shall find that computability issues will eventually be found to have a deep relevance to future physical theory, but only very little use of these ideas has so far been made in mathematical physics [^1] By computability issues Penrose alluded to undecidability, as he referred to Gödel and Turing in previous paragraphs, and provided as examples of this very little use the results of Komar (1964) and Geroch and Hartle (1986) that will be commented on in chapter 4..
Sometimes simple ideas do not have simple implementations, and taking undecidability to the realm of natural phenomena described by physical laws did not turn out to be easy at all. It is the goal of this review to show that the words of Bennett and Penrose might have to be reconsidered, since in the last decade an increasing number of central problems in physics, in particular many motivated by quantum information theory, have been proven undecidable.
@@
    Input: the description of an algorithm. Output: Yes or No, depending on whether the algorithm will halt or loop forever.
@@
A decision problem is called undecidable if no algorithm solves it. Computational problems without algorithms are called uncomputable. It is important to notice here that this is not a question of efficiency. That is, we are not asking for an algorithm whose runtime is efficient (e.g. polynomial) with respect to the number of bits of the input. If a problem is undecidable or uncomputable, there exists no algorithm for it, no matter how inefficient. Problem (c) above, known as the halting problem, was the first undecidable problem, discovered by Turing.
@@
In his famous first incompleteness theorem, Gödel proved that any reasonable set of axioms will always leave some statements that are axiomatically independent.
@@
Since mathematics describes physical systems with such unreasonable effectiveness —as E.P. Wigner wrote in his famous essay (Wigner (1960))— we might think that we can directly export to physics the undecidability that lives in mathematics. It is important to clarify that undecidability is not a feature of the physical system; it is a feature of the mathematical model we use to describe that physical system. It might not be possible to prove that a mathematical model of a physical system has a certain property, or that the property takes a certain concrete value. Yet that property might be experimentally measurable, and when it is measured on the real physical system, some concrete value must necessarily be obtained. However, this does not imply that physical systems can solve problems that are “beyond” mathematical reasoning. It may simply imply that mathematical models of physical systems are often idealizations that cannot be fully attained by real physical systems. Since any finite problem is decidable simply by enumerating the solutions, an undecidable problem must necessarily conceal an infinite somewhere: infinitely many instances, infinitely many particles, infinite precision. None of these idealized limits are directly accessible experimentally. However, as we will discuss further in section 6.2, undecidability in some idealized mathematical limit of a physical system can be reflected in surprising – sometimes even qualitatively new – (but still in principle computable) physical properties in the real, effectively finite physical system.
@@ 1.1 Complementary readings
K. Svozil has explored for many years the consequences in physics and philosophy of Gödel’s and Turing’s works (including as well most of the other mathematical systems discussed here) in various of his articles and books with plenty of references (Svozil (1994, 1995, 1997, 2005, 2011, 2018)).
The book by G. Chaitin, G. Doria and N. da Costa (Chaitin et al. (2011)) is an excellent informal introduction to undecidability in mathematics. It starts with Gödel and Turing, focuses on algorithmic information theory, with consequences in physics (including some of the results reviewed here), and with a more speculative final chapter on open problems and future developments.
C. Bennett wrote a number of interesting articles explaining some results about undecidability in physics that are covered as well in this review (Bennett (1990, 1995)). J.D. Barrow provides interesting “informal thoughts” about the implications of Gödel’s theorems in physics (Barrow (2000, 2011)), wondering whether all their assumptions can be applied to physical theories. There are various reviews treating undecidability in physics from the point of view of algorithmic information theory, such as the articles by C. Calude (Calude et al. (1995); Calude (2002a, 2005); Calude and Stay (2007)). One can also find in Cooper and Soskova (2017) a recent essay by various authors about the relevance of incomputability for the real world in different disciplines including physics.
Finally, there are plenty of references about undecidability in mathematics, like Davis (1977). More recent reviews are Margenstern (2000); Poonen (2014) covering undecidability and universality in different mathematical systems including Turing Machines, Diophantine equations, word problems, Post systems, molecular computations, register machines, neural networks, cellular automata and tilings. Goodman-Strauss (2010) covers most of these systems in a more informal and didactic way.
@@
Euclid’s Elements laid the foundations of mathematics in Ancient Greece and his axiomatic system remained untouched for more than two millennia. But the famous fifth postulate of geometry (the parallel postulate) was always under scrutiny [^2] Indeed, the proof of the independence of this postulate from the remaining axioms of plane geometry can be considered the first example of a mathematical statement which is proved undecidable with respect to an accepted axiomatic system (da Costa and Doria (1991b)).. In the 19th century mathematicians discovered that the parallel postulate could be changed or dropped entirely, leading to non-euclidean geometries. Besides opening a new fascinating world of curved geometries, mathematicians began to think again about the foundations of mathematics, and about the possibility of establishing a definite system of axioms and rules of inference powerful enough to construct the edifice of mathematics with absolute rigor.
The revolution in the foundations of mathematics that would come in the early 20th century was initiated in the late 19th century by G. Cantor, with his work on set theory and—to the great disturbance of the mathematical establishment of the epoch—with his infinite hierarchy of infinities. Practically ignored by his contemporaries, his goal of axiomatizing mathematics based on set theory was continued by G. Frege. After many years of hard work, Frege finally presented his first book in 1893. While the second one was being printed in 1903, he received a letter from B. Russell pointing to an insuperable contradiction in his system, now known as Russell’s Paradox: The set of all sets that are not members of themselves [^3] Frege immediately admitted the inconsistency in his system and never recovered from it. Russell wrote: “As I think about acts of integrity and grace, I realize that there is nothing in my knowledge to compare with Frege’s dedication to truth. His entire life’s work was on the verge of completion, much of his work had been ignored to the benefit of men infinitely less capable, his second volume was about to be published, and upon finding that his fundamental assumption was in error, he responded with intellectual pleasure clearly submerging any feelings of personal disappointment. It was almost superhuman and a telling indication of that of which men are capable if their dedication is to creative work and knowledge instead of cruder efforts to dominate and be known” (Irvine and Deutsch (2021)) (Irvine and Deutsch (2021)).
@@ 2.1 Gödel
Using prime numbers, Gödel managed to construct an ingenious coding system —the Gödel numbering— to assign a number to any symbol in a system F, in such a way that any group of symbols —a sentence— is coded by a unique number. With this powerful tool, he managed to encode into a number one particular sentence G that refers to itself: G is not provable in the system F [^6] In the original article, Gödel cited the analogy of his result with the liar paradox This sentence is false (Gödel (1931))
@@
At the time, Gödel’s work was widely ignored by his colleagues except by John von Neumann who understood immediately its tremendous impact, later worked on it, and wrote a letter to Gödel announcing that he had obtained an important corollary. Gödel replied that that result had been already sent for publication by himself: the Second Incompleteness Theorem, which proved that for any consistent system including elementary arithmetic, the consistency of the system cannot be proved in itself.
@@ 2.2 Church and Turing
In tackling this question, the first important step was to precisely define algorithm, or the recipe to calculate. Two people arrived at two different solutions: Alonzo Church with his Lambda-calculus (Church (1936)) and Alan Turing with his Turing Machine (Turing (1936)). Both ways of computing were later shown to be equivalent, and they served to relate the abstract mathematical concept of computation with concrete, definable processes via the Church-Turing Thesis.
@@
Even Gödel was quite convinced of this correspondence. He wrote: “The resulting definition of the concept of mechanical by the sharp concept of “performable by a Turing machine” is both correct and unique. […] Moreover it is absolutely impossible that anybody who understands the question and knows Turing’s definition should decide for a different concept.” (Copeland and Shagrir (2019)).
Since its formulation in the 1930’s there have been many attempts to prove or refute this thesis, as well as discussions of the status of this “thesis” and whether or not it is amenable to proof at all. All without any definitive resolution. It remains a thesis, not a theorem, but one widely accepted to be true by most mathematicians and computer scientists. We refer to Svozil (1994, 1997); Nielsen (2004); Davis (2006); Chaitin et al. (2011); Copeland and Shagrir (2019) for extensive discussions about the Church-Turing Thesis and subsequent extensions and variants.
@@
Both researchers found a negative answer to the decision problem, but Turing’s solution seemed more related to real-world problems. In fact, he considered the following ordinary problem:
@@
Input: the description of an algorithm. Output: YES or NO depending whether the algorithm halts or not.
@@ 2.3 Rice
In 1953, H.G. Rice proved a strong extension of Turing halting problem for computer science (Rice (1953)). Turing proved that in general it is impossible to know whether a program (Turing Machine) will halt or not. With a conceptually relatively simple proof, Rice showed that (informally) in general it is impossible to know what a program does to its input (see Hopcroft et al. (2006), Chaitin et al. (2011) for a formal statement of the theorem). If you are given an arbitrary program the only way to know what it does is running it. For instance, the questions:
@@ 2.4 Algorithmic Information Theory
Soon after that, a new incompleteness theorem was proven by G. Chaitin, in connection with ‘Algorithmic information theory (AIT)’; a theory founded independently by Solomonoff (1964), Kolmogorov (1965) and Chaitin (1966) in the 1960s, that revolves around the notion of complexity.
They managed to establish a precise and formal definition of a previously vague concept using information theory and computer science [^7] According to Chaitin, “AIT is the result of putting Shannon’s information theory and Turing’s computability theory into a cocktail shaker and shaking vigorously” (Calude (2002b)).. Broadly speaking, the Kolmogorov complexity [^8] Also known as Solomonoff-Kolmogorov-Chaitin complexity. of a string of bits is defined as the size in bits of the smallest computer program – the elegant program– that produces that string as output.
@@
By approaching incompleteness from the perspective of AIT, Chaitin showed Gödel undecidability to be pervasive. In an informal way: with a small amount of information (your formal axiomatic system) you can’t prove things that contain more information …and there are infinitely many things to prove. In Chaitin’s words:
With Gödel it looks surprising that you have incompleteness, that no finite set of axioms can contain all of mathematical truth. With Turing incompleteness seems much more natural. But with my approach, when you look at program size, I would say that it looks inevitable. Wherever you turn, you smash up against a stone wall and incompleteness hits you in the face! (Chaitin and Davies (2007)).
Three results about undecidability in quantum measurements that are based on AIT are commented on in sections 5.1.2 and 5.2.1 (Brukner (2009); Paterek et al. (2010); Trejo et al. (2023); Purcell et al. (2024)). General references about AIT and Kolmogorov complexity can be found in Chaitin (2003), Calude (2002b), Li and Vitnyi (2008), Chaitin et al. (2011).
@@ 3 Undecidabilty in Mathematics
Since all those pioneering works, many other problems in several areas of mathematics and computer science have shown to be undecidable. We will review those which are relevant for this review in section 3.4. But first, we need to first define the notion of undecidability more rigorously.
@@ 3.1 Gödel undecidability: axiomatic independence
Any consistent [^10] Consistent means that there is no sentence such that that sentence and its negation are both provable. formal system F within which a certain amount of elementary arithmetic [^11] See Ernest Nagel (2001); Franzén (2005); Raatikainen (2018) to clarify what certain amount of elementary arithmetic means in both theorems. In short, anything at least as strong as Robinson arithmetic, a first-order logic theory with seven axioms, is sufficient. Robinson arithmetic is a fragment of the standard axiomatisation of arithmetic via the Peano axioms. can be carried out is incomplete; i.e. there are statements in the language of F which can neither be proved nor disproved in F (Raatikainen (2018)).
@@
For any consistent formal system F within which a certain amount of elementary arithmetic11 can be carried out, the consistency of F cannot be proven in F itself (Raatikainen (2018)).
@@
As famous historical examples of axiomatic independence, we can cite Euclid’s fifth axiom (see footnote 2), the continuum hypothesis, and the axiom of choice in set theory (see for instance Cohen (2008)).
In this review we are not going to delve deeper into Gödel’s theorems and their proofs, since undecidability results in physics are mostly based on Turing’s undecidability (see Fig. 1). For rigorous and didactic references on Gödel’s results see Ernest Nagel (2001); Franzén (2005); Smith (2007); Raatikainen (2018).

### v2

@@ Abstract
The study of undecidability in problems arising from physics has experienced a renewed interest, mainly in connection with quantum information problems. The goal of this review is to survey this recent development. After a historical introduction, we first explain the necessary results about undecidability in mathematics and computer science. Then we briefly review the first results about undecidability in physics which emerged mostly in the 80s and early 90s. Finally, we focus on the most recent contributions, which we divide in two main categories: many body systems and quantum information problems.
journal: Physics Reports
[…]
[ALC]organization=Dpto. Automática, Escuela Politécnica, Universidad de Alcalá, postcode=28805, state=Madrid, country=Spain
[…]
[UCL]organization=Department of Computer Science, UCL, city=London, country=UK
[…]
[NTU]organization=Centre for Quantum Technologies, National University of Singapore, addressline=3 Science Drive 2, postcode=117543, country=Singapore
[…]
[CQT]organization=Nanyang Quantum Hub, School of Physical and Mathematical Sciences, Nanyang Technological University, postcode=639673, country=Singapore
[…]
@@ 1 Introduction
In a note about one of those attempts, C. Bennett wrote ([1]):
Alluding to the effects on twentieth-century physics of quantum mechanics and general relativity, the cosmologist John Wheleer once remarked that the world is simpler and stranger than we can imagine. An analogous dose of strange simplicity was delivered to mathematics in the 1930s by the discoveries of Gödel, Turing and their contemporaries in the theory of proof and computation. Yet, there have been surprisingly few applications of these ideas to physics. In particular, the phenomenon of ’undecidability’ has not asserted itself in a simple, uncontrived physical setting the way that deterministic chaos, a far weaker property, has.
And some years later, R. Penrose commented ([2] p378):
It is perhaps remarkable, in view of the close relationship between mathematics and physics, that issues of such basic importance in mathematics as transfinite set theory and computability have as yet had a very limited impact on our description of the physical world. It is my own personal opinion that we shall find that computability issues will eventually be found to have a deep relevance to future physical theory, but only very little use of these ideas has so far been made in mathematical physics [^1] By computability issues Penrose alluded to undecidability, as he referred to Gödel and Turing in previous paragraphs, and provided as examples of this very little use the results of [3] and [4] that will be commented on in chapter 4..
Sometimes simple ideas do not have simple implementations, and taking undecidability to the realm of natural phenomena described by physical laws did not turn out to be easy at all. One of the goals of this review is to show that the situation has changed since the words of Bennett and Penrose were written, as in the last decade an increasing number of central problems in physics (many of them motivated by quantum information theory), have been proven undecidable.
@@
    Input: the description of an algorithm and an input to it. Output: Yes or No, depending on whether the algorithm will halt or loop forever on the given input.
@@
A decision problem is called undecidable if no algorithm solves it. Computational problems without algorithms are called uncomputable. It is important to notice here that this is not a question of efficiency. That is, we are not asking for an algorithm whose runtime is efficient (e.g. polynomial) with respect to the number of bits of the input. If a problem is undecidable or uncomputable, there exists no algorithm for it, no matter how inefficient. Problem (c) above, known as the halting problem, was the first undecidable problem, discovered independently by A. Church and A. Turing in 1936.
@@
In his famous first incompleteness theorem, Gödel proved that any reasonably powerful set of axioms will always leave some statements that are axiomatically independent.
@@
Since mathematics describes physical systems with such unreasonable effectiveness —as E.P. Wigner wrote in his famous essay [5]— we might think that we can directly export to physics the undecidability that lives in mathematics. It is important to clarify that undecidability is not a feature of the physical system; it is a feature of the mathematical model we use to describe that physical system. It might not be possible to prove that a mathematical model of a physical system has a certain property, or that the property takes a certain concrete value. Yet that property might be experimentally measurable, and when it is measured on the real physical system, some concrete value must necessarily be obtained. However, this does not imply that physical systems can solve problems that are “beyond” mathematical reasoning. It may simply imply that mathematical models of physical systems are often idealizations that cannot be fully attained by real physical systems. Since any finite problem is decidable simply by enumerating the solutions, an undecidable problem must necessarily conceal an infinite somewhere: infinitely many instances, infinitely many particles, infinite precision. None of these idealized limits are directly accessible experimentally. However, as we will discuss further in section 6.2, undecidability in some idealized mathematical limit of a physical system can be reflected in surprising – sometimes even qualitatively new – (but still in principle computable) physical properties in the real, effectively finite physical system.
@@ 1.1 Complementary readings
K. Svozil has explored for many years the consequences in physics and philosophy of Gödel’s and Turing’s works (including as well most of the other mathematical systems discussed here) in various of his articles and books with plenty of references [6, 7, 8, 9, 10, 11].
The book by G. Chaitin, G. Doria and N. da Costa [12] is an excellent informal introduction to undecidability in mathematics. It starts with Gödel and Turing, focuses on algorithmic information theory, with consequences in physics (including some of the results reviewed here), and with a more speculative final chapter on open problems and future developments.
C. Bennett wrote a number of interesting articles explaining some results about undecidability in physics that are covered as well in this review [1, 13]. J.D. Barrow provides interesting “informal thoughts” about the implications of Gödel’s theorems in physics, wondering whether all their assumptions can be applied to physical theories [14, 15]. There are various reviews treating undecidability in physics from the point of view of algorithmic information theory, such as the articles by C. Calude [16, 17, 18, 19]. One can also find in [20], edited by Cooper and Soskova, a recent essay by various authors about the relevance of incomputability for the real world in different disciplines including physics.
Finally, there are plenty of references about undecidability in mathematics, like the one of Davis [21] or, more recently, the ones by Margenstern [22] and Poonen [23]. They cover undecidability and universality in different mathematical systems including Turing Machines, Diophantine equations, word problems, Post systems, molecular computations, register machines, neural networks, cellular automata and tilings. Goodman-Strauss covers in [24] most of these systems in a more informal and didactic way.
@@
Euclid’s Elements laid the foundations of mathematics in Ancient Greece and his axiomatic system remained untouched for more than two millennia. But the famous fifth postulate of geometry (the parallel postulate) was always under scrutiny [^2] Indeed, the proof of the independence of this postulate from the remaining axioms of plane geometry can be considered the first example of a mathematical statement which is proved undecidable with respect to an accepted axiomatic system [25].. In the 19th century mathematicians discovered that the parallel postulate could be changed or dropped entirely, leading to non-euclidean geometries. Besides opening a new fascinating world of curved geometries, mathematicians began to think again about the foundations of mathematics, and about the possibility of establishing a definite system of axioms and rules of inference powerful enough to construct the edifice of mathematics with absolute rigor.
The revolution in the foundations of mathematics that would come in the early 20th century was initiated in the late 19th century by G. Cantor, with his work on set theory and—to the great disturbance of the mathematical establishment of the epoch—with his infinite hierarchy of infinities. Practically ignored by his contemporaries, his goal of axiomatizing mathematics based on set theory was continued by G. Frege. After many years of hard work, Frege finally presented his first book in 1893. While the second one was being printed in 1903, he received a letter from B. Russell pointing to an insuperable contradiction in his system, now known as Russell’s Paradox: The set of all sets that are not members of themselves [^3] Frege immediately admitted the inconsistency in his system and never recovered from it. Russell wrote: “As I think about acts of integrity and grace, I realize that there is nothing in my knowledge to compare with Frege’s dedication to truth. His entire life’s work was on the verge of completion, much of his work had been ignored to the benefit of men infinitely less capable, his second volume was about to be published, and upon finding that his fundamental assumption was in error, he responded with intellectual pleasure clearly submerging any feelings of personal disappointment. It was almost superhuman and a telling indication of that of which men are capable if their dedication is to creative work and knowledge instead of cruder efforts to dominate and be known” [26] [26].
@@ 2.1 Gödel
Using prime numbers, Gödel managed to construct an ingenious coding system —the Gödel numbering— to assign a number to any symbol in a system F, in such a way that any group of symbols —a sentence— is coded by a unique number. With this powerful tool, he managed to encode into a number one particular sentence G that refers to itself: G is not provable in the system F.  [^6] In the original article, Gödel cited the analogy of his result with the liar paradox This sentence is false [27]
@@
John von Neumann attended Gödel’s conference in Königsberg and understood immediately its tremendous impact; later worked on it; and wrote a letter to Gödel announcing that he had obtained an important corollary. Gödel replied that the result had been already sent for publication by himself [28]: the Second Incompleteness Theorem, which proved that for any consistent system including elementary arithmetic, the consistency of the system cannot be proved in itself.
@@ 2.2 Church and Turing
In tackling this question, the first important step was to precisely define algorithm, or the recipe to calculate. Two people arrived at two different solutions: Alonzo Church with his Lambda-calculus [29] and Alan Turing with his Turing Machine [30]. Both ways of computing were later proved to be equivalent by Church, Turing and S. Kleene [31], and they served to relate the abstract mathematical concept of computation with concrete, definable processes via the Church-Turing Thesis.
@@
Even Gödel was quite convinced of this correspondence. He wrote: “The resulting definition of the concept of mechanical by the sharp concept of “performable by a Turing machine” is both correct and unique. […] Moreover it is absolutely impossible that anybody who understands the question and knows Turing’s definition should decide for a different concept” [32].
Since its formulation in the 1930’s there have been many attempts to prove or refute this thesis, as well as discussions of the status of this “thesis” and whether or not it is amenable to proof at all. No definitive resolution has been reached. It remains a thesis, not a theorem, but one widely accepted to be true by most mathematicians and computer scientists. We refer to [6, 8, 33, 34, 12, 32, 31] for extensive discussions about the Church-Turing Thesis and subsequent extensions and variants.
@@
Turing continued to investigate the real-world implications of his invention and considered the following ordinary problem
@@
Input: the description of an algorithm.
Output: YES or NO depending whether the algorithm halts or not.
@@ 2.3 Rice
In 1953, H.G. Rice proved a strong extension of Turing halting problem for computer science [35]. Turing proved that in general it is impossible to know whether a program (Turing Machine) will halt or not. With a conceptually relatively simple proof, Rice showed that (informally) in general it is impossible to know what a program does to its input (see [36], [12] for a formal statement of the theorem). If you are given an arbitrary program the only way to know what it does is running it. For instance, the questions:
@@ 2.4 Algorithmic Information Theory
Soon after that, a new incompleteness theorem was proven by G. Chaitin, in connection with ‘Algorithmic information theory (AIT)’; a theory founded independently by R. Solomonoff [37], A. Kolmogorov [38] and G. Chaitin [39] in the 1960s, that revolves around the notion of complexity.
They managed to establish a precise and formal definition of a previously vague concept using information theory and computer science [^7] According to Chaitin, “AIT is the result of putting Shannon’s information theory and Turing’s computability theory into a cocktail shaker and shaking vigorously” [40].. Broadly speaking, the Kolmogorov complexity [^8] Also known as Solomonoff-Kolmogorov-Chaitin complexity. of a string of bits is defined as the size in bits of the smallest computer program – the elegant program– that produces that string as output.
@@
By approaching incompleteness from the perspective of AIT, Chaitin showed Gödel undecidability to be pervasive. In an informal way: with a small amount of information (your formal axiomatic system) you cannot prove things that contain more information …and there are infinitely many things to prove. In Chaitin’s words [41]:
With Gödel it looks surprising that you have incompleteness, that no finite set of axioms can contain all of mathematical truth. With Turing incompleteness seems much more natural. But with my approach, when you look at program size, I would say that it looks inevitable. Wherever you turn, you smash up against a stone wall and incompleteness hits you in the face!
Four results about undecidability in quantum physics that are based on AIT are commented on in sections 5.1.2, 5.2.1 ((Brukner (2009) [42]; Paterek et al. (2010) [43]; Trejo et al. (2023) [44]; Purcell et al. (2024) [45]). General references about AIT and Kolmogorov complexity can be found in [46], [40], [47], [12].
@@ 3 Undecidabilty in Mathematics
Since all those pioneering works, many other problems in several areas of mathematics and computer science have shown to be undecidable. We will review those which are relevant for this review in section 3.4. But first, we need to define the notion of undecidability more rigorously.
@@ 3.1 Gödel undecidability: axiomatic independence
Any consistent [^10] Consistent means that there is no sentence such that the sentence and its negation are both provable. formal system F within which a certain amount of elementary arithmetic [^11] See [48, 49, 28] to clarify what certain amount of elementary arithmetic means in both theorems. In short, anything at least as strong as Robinson arithmetic, a first-order logic theory with seven axioms, is sufficient. Robinson arithmetic is a fragment of the standard axiomatization of arithmetic via the Peano axioms. can be carried out is incomplete; i.e. there are statements in the language of F which can neither be proved nor disproved in F [28].
@@
For any consistent formal system F within which a certain amount of elementary arithmetic11 can be carried out, the consistency of F cannot be proven in F itself [28].
@@
As famous historical examples of axiomatic independence, we can cite Euclid’s fifth axiom (see footnote 2), the continuum hypothesis, and the axiom of choice in set theory (see for instance [50]).
In this review we are not going to delve deeper into Gödel’s theorems and their proofs, since undecidability results in physics mostly derive from Turing’s halting problem. For rigorous and didactic references on Gödel’s results see [48, 49, 51, 28].